OpenStudy (shesolitt):

I WILL MEDAL AND FAN!!! Choose one of the factors of 1331x3 − 8y3. 4 11x − 2y <<

OpenStudy (shesolitt):

@RhondaSommer @jabez177

OpenStudy (shesolitt):

I'm actually not sure if its B or D

OpenStudy (latinc):

1331x3-8y^3 / 11x-2y if you don't get a remainder its a factor

OpenStudy (shesolitt):

so I was correct? the answers B? @latinC

OpenStudy (latinc):

Yes:)

OpenStudy (shesolitt):

THANK YOU SO MUCH!

OpenStudy (anonymous):

\[\left( 1331x^3-8y^3 \right)=\left( 11x \right)^3-\left( 2y \right)^3\] \[a^3-b^3=\left( a-b \right)\left( a^2+ab+b^2 \right)\]

OpenStudy (studio-games):

Please do not give direct answers @latinC